OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (6, 2) and (4, 8)? y = x +

OpenStudy (leader):

slope= 6/-2 = -3 So y=-3x+b 2=-3(6)+b 2=-18+b b=2 So y=-3x+20

OpenStudy (leader):

oops b=20

OpenStudy (anonymous):

thnx